Calculate the concentration of copper(II) chloride in solution, in mol/L, if 125 g of CuCl2 crystals are dissolved in 0.50 L of water. Then, calculate the chloride ion concentration in mol/l.


1
Expert's answer
2020-09-24T08:14:53-0400

molarmass of copper(II)chloride is 134g/mol, weight of dissolve cucl2 in solution is 125g,

no of moles of cucl2 dissolve in solution =weight of cucl2 dissolve/molarmass of cucl2

= 125/134=0.93 mol,

0.93 mol cucl2 dissolve in 0.50 L solution,hence the concentration of the solution in mol/L is (0.93/0.5)=1.86,

one mole cucl2 under dissociation produce 2mol chloride ions,hence the chloride ions concentration is (1.86*2)=3.72mol/L